# Low-Light Video Enhancement via Spatial-Temporal Consistent Decomposition

### v2

@@ 4.6 The Effects of CFIM Structure
RetinexFormer 26.56 0.79 22.80 0.77 29.15 0.82 25.40 0.89
RetinexFormer+CFIM 27.42 0.80 23.75 0.79 29.87 0.83 26.38 0.90
Table 5: The experiments to show the effects of CFIM.
@@ 4.7 Robustness Towards Correspondences
In this section, we explore the robustness of our framework with respect to incorrect and insufficient correspondences. In the first experiment, we introduce random noise perturbations to the computed correspondences, modifying the location values of the correspondences (the perturbation range is sampled from $-20\sim 20$ pixels, which falls within the normal error range of current correspondence estimation methods (Edstedt et al., 2023)). As shown in Table 6, the results with perturbed correspondences (denoted as Ours with per.) show a performance drop compared to the unperturbed case. However, the performance remains superior to that of most SOTA baselines, indicating that our framework is robust to a certain of erroneous correspondences, which are likely to occur in areas with complex textures or other challenging scenarios. In the second experiment, we address the issue of insufficient correspondences, which can arise, for instance, in regions with occlusions. To simulate this, we randomly reduce the number of correspondences to 10% of the originally detected ones. The results (Ours with reduced) are presented in Table 6, where we observe that our method continues to outperform almost all baselines in Table 1, albeit with a smaller margin than the original setting. This suggests that even with fewer correspondences, the remaining ones still provide valuable information and they are in accord with the other constraints. This further demonstrates the robustness of our method.
@@
Table 6: The experiments to show the robustness of our framework towards correspondences.

### v3

@@ 4.6 The Effects of CFIM Structure
R.F. 26.56 0.79 22.80 0.77 29.15 0.82 25.40 0.89
R.F.+CFIM 27.42 0.80 23.75 0.79 29.87 0.83 26.38 0.90
Table 5: The effects of CFIM. “R.F." denotes RetinexFormer.
@@ 4.7 Robustness Towards Correspondences
In this section, we explore the robustness of our framework with respect to incorrect and insufficient correspondences. In the first experiment, we introduce random noise perturbations to the computed correspondences, modifying the location values of the correspondences (the perturbation range is sampled from $-20\sim 20$ pixels, which falls within the normal error range of current correspondence estimation methods Edstedt et al. (2023)). As shown in Table 6, the results with perturbed correspondences (denoted as “Ours with per.") show a performance drop compared to the unperturbed case. However, the performance remains superior to that of most SOTA baselines, indicating that our framework is robust to a certain of erroneous correspondences, which are likely to occur in areas with complex textures or other challenging scenarios. In the second experiment, we address the issue of insufficient correspondences, which can arise, for instance, in regions with occlusions. To simulate this, we randomly reduce the number of correspondences to 10% of the originally detected ones. The results (“Ours with red.") are presented in Table 6, where we observe that our method continues to outperform almost all baselines in Table 1, albeit with a smaller margin than the original setting. This suggests that even with fewer correspondences, the remaining ones still provide valuable information and they are in accord with the other constraints. This further demonstrates the robustness of our method.
@@
Table 6: The robustness of our framework towards correspondences.